OpenStudy (anonymous):

(x-2)^+(x-2)-12=0 solve using an appropriate technique????

OpenStudy (anonymous):

x^2-4x+4+x-2-12 =x^2-3x-10=0 =x^2-5x+2x-10=0 (x-5)(x+2)=0 x=5 x=-2

OpenStudy (anonymous):

sai1234 you are awesome thank you.

OpenStudy (anonymous):

(x-2)^2 +(x-2) -  12 =0 Let y= x-2 substituting we get Y ^2 + Y -12 = 0 Y^2 + 4 Y - 3Y -12 = 0 Y( Y + 4) -3(Y+ 4)= 0 (Y+ 4)(Y-3)=0 x-2+4= 0, x-2-3=5 X=-2, x= 10

OpenStudy (anonymous):

Sorry it is x-2-3=0 X= 5

OpenStudy (anonymous):

This is the substituting technique to solve it in a simple manner